OpenStudy (anonymous):

Let f(x) = 7x – 13. Find f–1(x).

OpenStudy (anonymous):

x+3/7

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

replace the function notation with y y = 7x – 13 exchange the positions of the x and y variables x = 7y - 13 then solve for y x+13 = 7y dividing by 7\[y=\frac{x+13}{7}\]or\[f ^{-1}(x)=\frac{1}{7}x-\frac{13}{7}\]

OpenStudy (anonymous):

Note: I made a minus sign error on my last line; it is supposed to be +13/7 (se it is correct in the second to last algebraic line). correct is\[f ^{-1}(x)=\frac{1}{7}x+\frac{13}{7}\]

OpenStudy (anonymous):

You are aware that inverse functions are symmetric about the line y=x, here is that relationship